OpenStudy (isbella):

test help!1)Set up a double integral for calculating the flux of F⃗ =3xi⃗ +yj⃗ +zk⃗ through the part of the surface z=−5x−2y+4 above the triangle in the xy-plane with vertices (0,0), (0,3), and (2,0), oriented upward. Flux = ∫BA∫DC A = B = C = D = Flux = ∬SF⃗ ⋅dA⃗ =

OpenStudy (anonymous):

Are you familiar with Electricity and Magnetism Physics?

OpenStudy (anonymous):

First let's think of the definition of flux, it is noted by \[\phi _{0} = EA cos(\theta)\]

OpenStudy (anonymous):

You find dA for the the plane, you find \(dA= \sqrt {30} \)

OpenStudy (anonymous):

You find the normal upward vector to the plane (5,2,1)

OpenStudy (anonymous):

You find (5,2,1) dot ( 3 x , y, -5 x - 2 y + 4)=4 + 10 x

OpenStudy (anonymous):

\[ \int_0^2 \left(\int_0^{3-\frac{3 x}{2}} \sqrt{30} (10 x+4) \, dy\right) \, dx=32 \sqrt{30} \]

OpenStudy (anonymous):

The bound are for the triangle in the xy plane